OpenStudy (anonymous):

The sixth grade class has 325 students; the ratio was 24 boys to 26 girls. Out of the 325 students, how many boys are there? please explain to me how you do this!!

OpenStudy (anonymous):

24:26 boys is the ratio

OpenStudy (anonymous):

not the ratio! how many boys there are!

OpenStudy (anonymous):

therefore if u simplify that ratio by dividing by 2 ull get 12:13. Now the total no.of boys is 12/13+12*325 =12/25*325=156boys

OpenStudy (anonymous):

u understand

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i got 168

OpenStudy (anonymous):

well thats my best only answer as far as my experience

OpenStudy (s):

x = boys 26x/24 = girls So, x + 26x/24 = 325 24x/24 + 26x/24 = 325 50x/24 = 325 50x = 325 * 24 50x = 7800 x = 7800 / 50 x = 156 - boys 26x/24 = 26*156/24 = 169 - girls

OpenStudy (anonymous):

Set a proprtion 25 students - 12 boys 325 students - x boys \[x=\frac{325*12}{25}=156 boys \]

OpenStudy (anonymous):

thx! this really helps!

OpenStudy (anonymous):

You are welcome. Glad to help.